OpenStudy (anonymous):

x-y=11 2x+y=19 solve by Elimination

OpenStudy (anonymous):

Help me please

myininaya (myininaya):

Add the equations together

myininaya (myininaya):

2x+x-y+y=11+19

myininaya (myininaya):

3x+0=30 3x=30 x=10

myininaya (myininaya):

x-y=11 10-y=11 -y=11-10 -y=1 y=-1